A boat moves through the water with two forces acting on it. One is a 2.72×10 3 N forward push by the motor, and the other is a 1.76×10 3 N resistive force due to the water. What is the acceleration of the 1384.9 kg boat? Answer in units of m/s 2 If it starts from rest, how far will it move in 17.3 s? Answer in units of m
1
Expert's answer
2011-10-27T12:57:30-0400
Due to Newton's second law:
Fm - Fres = ma
where Fm is the force which
motor pushes and Fres is a resistive force.

a = (Fm - Fres)/m =
(2.72*10^3 - 1.76*10^3) / 1384.9 = 0.693 m/s^2.
The equation for displacement
is:
S = V0*t & + (at^2)/2, V0 = 0.
S = (at^2)/2 = (0.693*17.3^2)/2 = 103.7
m